# DiffClean: Diffusion-based Makeup Removal for Accurate Age Estimation

### v1

@@ Abstract
Accurate age verification can protect underage users from unauthorized access to online platforms and e-commerce sites that provide age-restricted services. However, accurate age estimation can be confounded by several factors, including facial makeup that can induce changes to alter perceived identity and age to fool both humans and machines. In this work, we propose DiffClean which erases makeup traces using a text-guided diffusion model to defend against makeup attacks. DiffClean improves age estimation (minor vs. adult accuracy by 4.8%) and face verification (TMR by 8.9% at FMR=0.01%) over competing baselines on digitally simulated and real makeup images.
1 Introduction
@@
Motivation. Our constant need to stay digitally connected has given rise to a wide range of online services, including social media platforms, chat rooms, and dating websites. Several online services rely on third-party selfie-based age verification services for user authorization (e.g., Yoti, Ondato, TrustStamp). But there have been reports of minors bypassing age verification checks exposing them to cyberbullying and online dangers. [^1] Children bypass age-checking measures on social media sites  [^2] Underage teens bypassing age restrictions on dating apps. In such scenarios, precise age estimation is critical in protecting underage users. Age estimation is intrinsically challenging due to complex genetic, environmental, demographic, and subjective factors. The problem is compounded when an individual puts on makeup. Facial makeup significantly alters the perception of facial features that serve as visual cues for age prediction [14, 41, 9]. The perceived age in younger people with makeup is generally overestimated, while, in older people, the predicted age is underestimated [52, 54]. Makeup can also fool automated age estimation frameworks [6]. Therefore, we need a reliable makeup-invariant age estimator.
Intuitively, we can improve the robustness of the age estimation algorithms to the presence of makeup by using “more data”. However, this immediately raises privacy concerns: we will need data on minors wearing makeup. The next strategy might be to generate synthetic makeup images for training age classifiers. But makeup can vary in terms of intensity, texture, color, region of application, and subjective preferences. So, synthetic data may not capture the full diversity of facial makeup. Thus, we propose to design a makeup removal framework to erase makeup traces digitally, and help restore the actual age. Additionally, our method also recovers identity cues confounded by makeup.
Our approach. Makeup removal can simultaneously benefit age estimation and face-based identity verification. To achieve this objective, we formulate makeup removal as an image translation task. Naturally, we adopt generative models that are adept at facial editing; specifically, seamless and controllable digital makeup transfer [30, 7, 24, 34, 56, 43, 48]. Our goal differs from previous work that focused on attribute editing and adversarial makeup transfer. We need to restore the age cues of original image that may have been perturbed after applying makeup. So, we propose DiffClean that uses a text-guided diffusion model to transform a face image with makeup to a face image without makeup using a combination of CLIP loss (for makeup detection), perceptual loss (for fidelity retention), biometric loss (for identity restoration), and age loss (for age restoration); see Fig. 1. DiffClean is efficient and can be plugged in as a pre-processing module to erase makeup traces, if present, for downstream facial analytics.
@@
  • We propose Diffclean, a novel text-guided diffusion-based makeup removal framework that digitally erases the traces of makeup present in face images, and improves age estimation (minor vs. adult accuracy by 4.8%) and face verification (TMR by 8.9% at FMR=0.01%) over state-of-the-art baselines.
  • We utilize text-guided CLIP loss for detecting makeup, perceptual loss for minimizing visual artifacts, and a combination of age and identity losses for restoring the facial cues affected by makeup. Our method shows improved generalizability over existing methods in terms of different makeup styles and demographic factors.
  • We demonstrate that our method works on both digitally simulated and real makeup images, thus capable of mitigating differential effects of makeup that may help protect underage users from cyber harms. We will release the codes for reproducible research.
@@
Figure 1: Overview of our method, DIFFCLEAN, that removes makeup on an example FFHQ image using text-guided diffusion model with a combination of CLIP loss, identity (ID) loss, Age loss and Perceptual (Visual) losses. Successful makeup removal results in shifting of softmax probabilities of $[p(withmakeup),p(withoutmakeup)]$ from $[0.76,0.24]$ to $[0.13,0.87]$, and reduces overestimation from 14.6 yrs to 10.4 yrs.
2 Related Work
@@
Facial Age Prediction. Classical approaches extract handcrafted features to design classification and regression models. [28] used statistical face models generated by applying Principal Component Analysis (PCA) on a set of face images for age estimation. Texture features such as Local Binary Pattern (LBP) and its variants have been used for reliable age estimation [17, 44, 59]. Biologically-inspired features have been used in [18, 20]. Deep-learning -based approaches rely on convolutional neural networks (CNNs) and transformers for improved robustness. A multi-stream CNN was used to learn high-dimensional structured information from patches in facial images for reliable age estimation in [58]. Deep CNN was used for age estimation in unconstrained settings in [38, 53, 29]. In [36], each age was treated as a rank to design an end-to-end learning approach to apply ordinal regression problem to age estimation using deep CNN. A multi-task learning approach for heterogeneous attribute estimation was adopted in [21]. Deep Expectation (DEX) [39] formulated age estimation by converting it into a classification-regression problem. SSRNet [57] adopted a compact soft stagewise regression model that employed a coarse-to-fine strategy to perform multi-class classification. Auxiliary facial cues were used to boost reliable age estimation in [33, 60]. Deep random forests were used for age estimation in [45]. FP-age [32] used face parsing attention module to incorporate facial semantics into the age estimation. Swin Transformer with attention-based convolution (ABC) for facial age estimation was proposed in [46]. In [27], a multi-input transformer-based model, MiVOLO, was designed for age and gender prediction using facial and full body information.
Figure 2: Makeup diversity due to variations in intensity (light vs. heavy), region of application (eye vs. lip), color (pink vs. red), texture (matte vs. gloss), etc. Example images are generated by synthetic retouching the original image from CelebA-HQ[25] using Google Gemini Flash 2.0.
@@
Age Estimation in Minors. While there is extensive research on age estimation for adults, limited work have focused on improving the age estimation performance for minor age groups. In [3], a finetuned VGG-based ‘children’ network was used for apparent age estimation of children of age from 0 to 12 years. [2] trained ResNet50 on underage facial dataset ‘VisAGe’ to improve age estimation for minors. The need for fast, accurate and robust age estimation and verification has gained traction recently. NIST Face Analysis Technology Evaluation (FATE) Age Estimation & Verification reports that subjects aged between 8 and 12 years tend to be overestimated to be 13 to 16 years old with false positives as high as $0.76$  [^3] NIST FATE Report; higher errors are reported for females than for males, specifically, in 17-year-olds [22]. Performance of facial age estimation algorithms can be affected by several factors like gender, ethnicity, image quality, facial expression, and makeup [1, 8].
@@
Impact of Cosmetic Makeup on Facial Analytics. [51] demonstrated the impact of facial makeup on automatic face recognition. [14] conducted a comprehensive study suggesting that makeup applied to different facial regions also has different effects on age prediction and further quantified the makeup effect as Young Index (YI). [13] observed that Identification Rate (IDR) decreases with increasing amount of makeup. Taking into account the adverse effects of makeup in face verification task, [5, 19, 23] focused on improving robustness of face recognition to makeup. Authors in [6] observed that facial cosmetics (both real makeup and synthetically retouched) affect automated age estimation causing a mean absolute error of up to 5.84 years. They further indicated that lip makeup had least impact on automated age estimation, followed by eye makeup, followed by full face makeup (eye + lip + foundation). Makeup has shown to produce differential effects in apparent (manual) age estimation. 20-year-old women appeared older when wearing full face makeup while 40- and 50-year-old women appeared younger when wearing makeup [10, 8, 41, 9]. Limited work discuss the influence of makeup on facial age estimation in minors [1, 2].
@@
Facial Makeup Transfer and Removal. Approaches for makeup transfer relied on GAN-based style-transfer networks to map reference makeup styles to target image [30, 7, 24, 34, 56, 50]. Later, diffusion models (and recently, multimodal LLMs) bridged this gap [62, 49, 40]; see examples of diverse makeup editing in Fig. 2. Makeup removal, on the other hand, poses a more challenging task where the goal is to reveal the face under makeup without adversely affecting the biometric and perceptual features. To assist with face verification, [31] integrated two adversarial networks, both at pixel and feature levels to generate identity-preserving synthetic non-makeup faces, while, [61] replaced convolution layers in GAN with resnet blocks to design makeup removal network. [40] proposed an one-for-all diffusion-based framework for a wide array of makeup tasks including makeup transfer and removal by modeling each as a domain translation problem. CLIP2Protect [43] and DiffAM [48] use adversarial makeup transfer for privacy with capability to generate ‘no makeup’ faces.
3 Proposed Method
@@
Makeup has an adverse effect on facial analytics, specifically, exposing minors to online harms by breaching automated age-based regulation checks. Therefore, addressing this issue in a principled way will have a positive societal impact. One workaround is to collect representative data of minors wearing makeup for improving robustness; this poses strong privacy concerns. So, we propose a generative model, Diffclean, that is designed to mitigate the age variations arising due to makeup. There are two challenges to designing this network: (1) we need a robust makeup style generator to synthesize data needed for training thus circumventing data collection; (2) we need a robust age estimator to serve as a proxy auxiliary classifier to guide DiffClean to restore age-specific cues.
@@
Figure 3: Synthetic makeup transfer results produced using EleGANt [56] on example images from UTKFace. The first column indicates the original image, the second column indicates the makeup style reference, and the third column indicates the images after makeup transfer.
For the makeup style generator, we used an existing network, EleGANt [56] that can perform locally-editable makeup transfer; see Fig. 3. Age estimation algorithms typically perform poorly on lower age groups which are most vulnerable to overestimation due to the application of makeup. Therefore, we first need an auxiliary age estimator that has been specifically fine-tuned on our target age groups (10-19 years). For the proxy age classifier, we used SSRNet [57] fine-tuned on UTKFace [63] dataset with weighted self-adjusting smoothed-$\mathcal{L}_{1}$ loss to aggressively penalize misclassifications in target age groups.
@@
DiffClean uses the ablated diffusion model (ADM) with guided diffusion from [12] with improved U-Net architecture containing increased attention heads at multiple resolutions, BigGAN residual stacks for upsampling and downsampling, deeper networks and rescaled residual connections; refer to  [12] for details. We used text prompts face with makeup and face without makeup for text-guided CLIP directional loss following [48]. The CLIP loss helps remove makeup traces conditioned on identity, age and perceptual constraints. For training the diffusion model, we sampled images from the UTKFace [63] dataset that consists of original face images and applied EleGANt to generate (w/o makeup, w/ makeup) pairs. We evaluated our trained DiffClean on FFHQ [26] dataset and used the exact age values provided by [11] in conjunction with fine-tuned SSRNet to ensure that the predicted age of the image after makeup removal is similar to the input image with makeup.
@@
Methodology. We describe the terminology and notation here. Original/Clean image (without makeup): $\textbf{I}_{o}$, Makeup image (with real or synthetic makeup), $\textbf{I}_{m}$, and Makeup removed image, $\textbf{I}_{g}$. Consider $\mathcal{M}$ as a real or synthetic makeup application operator, while $\mathcal{M}^{-1}$ represents the makeup removal operator. Thus, $\textbf{I}_{o}\xrightarrow[]{\mathcal{M}}\textbf{I}_{m}$, while $\textbf{I}_{m}\xrightarrow[]{\mathcal{M}^{-1}}\textbf{I}_{g}$. The goal is to model $\mathcal{M}^{-1}$ so that we can recover the clean image, $\textbf{I}_{g}\approx\textbf{I}_{o}$. Our proposed framework DiffClean approximates the role of $\mathcal{M}^{-1}$.
@@
Architecture: The makeup removal framework ($\mathcal{M}^{-1}$) used in this work is a generative model which is a pre-trained ADM [12] that adds noise to the input image (with makeup), $\textbf{I}_{m}$, to generate the noisy latent image during forward diffusion. The deterministic reverse diffusion through DDIM produces the makeup removed output, $\textbf{I}_{g}$ supervised by CLIP, biometric identity, age and perceptual losses.
Loss functions:
@@
Proxy Age Estimator: The age estimator, SSRNet is fine-tuned using a self-adjusting smoothed-$\mathcal{L}_{1}$ loss introduced in [15] and adaptively weighted for different age groups as follows:
@@
where,
$\mathcal{L}_{SL}(a,\hat{a})=\begin{cases}0.5\frac{(a-\hat{a}).^{2}}{\beta},&\text{if $a-\hat{a}<\beta$}\\ |a-\hat{a}|-0.5*\beta,&\text{otherwise}\end{cases}$
Here, $a$ is the ground truth age, $\hat{a}$ is the predicted age, and the parameter $\beta$ is adapted based on the running mean and variance. Note we ablated with different losses such as $\mathcal{L}_{1}$, $\mathcal{L}_{2}$ and Huber loss functions but empirically observed weighted self-adjusting smooth $\mathcal{L}_{1}$ loss to perform the best.
DiffClean: During training, we used four losses to fine-tune the diffusion model as follows.
CLIP-based Makeup Loss: The transition from makeup domain to non-makeup domain is guided via textual prompts ‘face without makeup’ and ‘face with makeup’ in CLIP space following [48]. The loss is one minus the cosine similarity between the image and text embeddings as follows:
@@
Biometric identity Loss: Our goal involves identity restoration of the makeup removed image with respect to the original image (without makeup). During training with synthetic makeup pairs, we have access to ($\textbf{I}_{o}$, $\textbf{I}_{m}$) to generate $\textbf{I}_{g}$. So, we compute weighted pairwise biometric identity losses as $\mathcal{L}_{og}$ and $\mathcal{L}_{mg}$. This is different than [48] that uses a single identity loss between input and output. The final identity loss is as follows:
@@
Perceptual Loss: To preserve the perceptual features, we use the perceptual loss $\mathcal{L}_{pips}$ with VGG [47] network and $\mathcal{L}_{1}$ loss between generated makeup-removed image $\textbf{I}_{g}$ and input image with makeup $\textbf{I}_{m}$.
Age Loss: Restoration of age cues requires explicit fine-tuning of the makeup removal module with age loss. The age loss is computed as the smoothed-$\mathcal{L}_{1}$ loss between the the ground truth age, $a_{i}$ of the original image, $\textbf{I}_{o}$, and the predicted age, $\hat{a_{i}}$ of the makeup removed image, $\textbf{I}_{g}$. We used the fine-tuned SSRNet model to predict $\hat{a_{i}}$. Note age loss needs a specific value, whereas, FFHQ-Aging provides age groups. So, we used the user metadata (age labels) provided by [11]. We introduce the SSRNet age loss as follows:
$\mathcal{L}_{\text{SSRNet-age}}=\frac{1}{n}\sum_{i=1}^{n}\mathcal{L}_{SL}(a_{i},\hat{a}_{i})$ (4)
[…]
@@
$\mathcal{L}_{total}=\lambda_{clip}\mathcal{L}_{clip}+\lambda_{id}\mathcal{L}_{id}+\lambda_{lpips}\mathcal{L}_{lpips}+\lambda_{L1}\mathcal{L}_{L1}+\lambda_{age}\mathcal{L}_{age}$ (6)
@@
where loss weights are tuned as $\lambda_{clip}=5.0$, $\lambda_{id}=1.0$, $\lambda_{lpips}=5.0$, $\lambda_{L1}=2.0$, and $\lambda_{age}=0.5$ for SSRNet age loss and $\lambda_{age}=5.0$ for CLIP age loss. Thus, we design two models: DiffClean with SSRNet age loss ($\mathcal{L}_{age}=\mathcal{L}_{\text{SSRNet-age}}$) and DiffClean with CLIP age loss ($\mathcal{L}_{age}=\mathcal{L}_{\text{CLIP-age}}$).
During testing, our model can accept any face image (w/o makeup or w/ real/synthetic makeup) as input. We perform age estimation using MiVOLO [27] (while training uses SSRNet) and MobileFace and FaceNet for identity verification (while training uses ArcFace IRSE50).
4 Experiments and Analysis
4.1 Datasets

### v2

@@ Abstract
Accurate age verification can protect underage users from unauthorized access to online platforms and e-commerce sites that provide age-restricted services. However, accurate age estimation can be confounded by several factors, including facial makeup that can induce changes to alter perceived identity and age to fool both humans and machines. In this work, we propose DiffClean which erases makeup traces using a text-guided diffusion model to defend against makeup attacks without requiring any reference image unlike prior work. DiffClean improves age estimation (minor vs. adult accuracy by 5.8%) and face verification (TMR by 5.1% at FMR=0.01%) compared to images with makeup. Our method is: (1) robust across digitally simulated and real-world makeup styles with high visual fidelity, (2) can be easily integrated as a pre-processing module in existing age and identity verification frameworks, and (3) advances the state-of-the art in terms of biometric and perceptual utility. Our codes are available at https://github.com/Ektagavas/DiffClean
Index Terms:
Makeup removal, Age estimation, Online safety
I Introduction
@@
Our constant need to stay digitally connected has given rise to a wide range of online services, including social media platforms, chat rooms, and dating websites. Several online services rely on digital age verification, typically, age verification from selfie-styled facial scans (e.g., Yoti, Ondato, TrustStamp). But there have been reports of minors bypassing age verification checks exposing them to cyberbullying and online dangers.  [^1] Children bypass age-checking measures on social media sites [^2] Underage teens bypassing age restrictions on dating apps. In such scenarios, precise age estimation is critical in protecting underage users. Recently, government agencies and commercial content providers are actively pushing for online facial age verification for protecting teenagers (US’s Identifying Minors Online, UK’s Online Safety Act, YouTube) [congress_2024, youtube_2025] from unauthorized and unmitigated access to adult-themed websites.
Figure 1: Overview of our method, DiffClean, that removes makeup on an example FFHQ image using text-guided diffusion model with a combination of CLIP loss, identity (ID) loss, Age loss and Perceptual (Visual) losses. Successful makeup removal results in shifting the softmax probabilities: [p(with makeup),p(without makeup)][p($\text${with makeup}),p($\text${without makeup})] from $[0.76,0.24]$ to $[0.13,0.87]$. Here, we use ViT-B/32 as the CLIP-based classifier. DiffClean successfully maintains biometric and perceptual quality while reducing the overestimated age from 14.6 yrs (with makeup) to 10.4 yrs (without makeup).
Our motivation. Age estimation is intrinsically challenging due to complex genetic, environmental, demographic, and subjective factors; this is compounded when an individual puts on makeup. Facial makeup significantly alters the perception of facial features affecting reliable age prediction [feng2012quantifying, russell2019differential, davis2022your]. The perceived age in younger people with makeup is generally overestimated, while in older people the predicted age is underestimated [voelkle2012let, watson2016you]. Makeup can be used to bypass automated age estimation [Chen2014ImpactOF], the primary issue addressed in this work, but we also look at its impact on face recognition [ueda2010influence]. Intuitively, we can design a makeup-invariant age estimator and face matcher, specifically designed for minors by using “more data”. However, this immediately raises privacy concerns: we will need data on minors wearing makeup. The next strategy might be to generate synthetic makeup images for training new models. But makeup can vary in terms of intensity, texture, color, region of application, and subjective preferences. So, synthetic data may not adequately capture the wide spectrum of variations possible in facial makeup.
Our goal. We propose to defend against makeup attacks by designing a makeup removal framework, known as DiffClean. It will be deployed as a plug-in module that will filter the selfie-styled facial input image to erase makeup traces, which will primarily benefit accurate age estimation, and associated secondary tasks, such as, identity verification.
Our approach. We formulate makeup removal as an image translation task, leveraging generative models known for seamless and controllable digital makeup transfer [li2018beautygan, chen2019beautyglow, jiang2020psgan, nguyen2021lipstick, yang2022elegant, shamshad2023clip2protect, sun2024diffam]. Unlike prior works that rely on attribute editing or reference-based adversarial transfer, our method DiffClean is a reference-free, text-guided diffusion model that transforms a face image with makeup to a face image without makeup using a combination of CLIP loss (for makeup detection), perceptual loss (for fidelity retention), biometric loss (for identity restoration), and age loss (for age restoration); see Fig. 1. Our method removes makeup traces if present in face images to primarily improve age estimation accuracy, and additionally, restore biometric cues to improve identity verification performance. Note that in absence of makeup, our method retains and does not degrade age and biometric markers.
@@
  1. 1.
    We propose Diffclean, a novel reference-free text-guided diffusion-based makeup removal framework that digitally erases the traces of makeup present in face images, and improves age estimation (minor vs. adult accuracy by 5.8%) and face verification (TMR by 5.1% at FMR=0.01%) compared to images with makeup, while retaining visual quality (SSIM=0.98, PSNR=35.64).
  2. 2.
    We utilize text-guided CLIP loss for detecting makeup, perceptual loss for minimizing visual artifacts, and a combination of age and identity losses for restoring the facial cues affected by makeup. The combined losses ensure that our model works on both makeup and non-makeup images without producing artifacts.
  3. 3.
    We empirically demonstrate that DiffClean works on both digitally simulated and real makeup images across variations in pose, illumination, facial accessories, and different makeup styles, while exhibiting demographic fairness. Our method can assist existing digital age verification platforms as a pre-processing module to protect underage users from cyber harms. Our method outperforms six (four GAN-based and two diffusion-based) baselines on four datasets. We have released the codes for reproducible research.
@@
II Related Work
@@
Facial Age Prediction.  [lanitis2004comparing] used statistical face models generated by applying Principal Component Analysis (PCA) on a set of face images for age estimation. Texture features such as Local Binary Pattern (LBP) and its variants have been used for reliable age estimation [LBP_2008, Localfeatures_2010, LBP_2012]. Biologically-inspired features have been used in [biofeatures_2009, HumansvsMC_2013]. Deep-learning -based approaches rely on convolutional neural networks (CNNs) and transformers for improved robustness. A multi-stream CNN was used to learn high-dimensional structured information from face patches in [yi2014age]. Deep CNN was used for age estimation in unconstrained settings in [ranjan2015unconstrained, wang2015deeply, levi2015age]. In [niu2016ordinal], each age was treated as a rank to apply ordinal regression problem to age estimation using deep CNN. A multi-task learning approach for heterogeneous attribute estimation was adopted in [han2017heterogeneous]. Deep Expectation (DEX) [rothe2018deep] formulated age estimation by converting it into a classification-regression problem. SSRNet [yang2018ssr] adopted a compact soft stagewise regression model to perform multi-class classification. Auxiliary facial cues were used to boost reliable age estimation in [liu2020facial, zeng2020soft]. Deep random forests were used for age estimation in [DeepForests_2021]. FP-age [lin2022fp] used face parsing attention module to incorporate facial semantics into age estimation. Swin Transformer with attention-based convolution (ABC) for facial age estimation was proposed in [Shi_2023]. In [kuprashevich2023MiVOLO], a multi-input transformer-based model, MiVOLO, predicted age and gender using facial and full body information.
@@
Age Estimation in Minors. Limited works have focused on improving age estimation performance in minors. In [antipov2016apparent], a finetuned VGG-based ‘children’ network was used for apparent age estimation of children of age from 0 to 12 years. [anda2020deepuage] trained ResNet50 on the underage facial dataset ‘VisAGe’ to improve the age estimation for minors. The need for fast, accurate, and robust age estimation and verification has gained traction recently. NIST Face Analysis Technology Evaluation (FATE) Age Estimation & Verification reports that subjects aged between 8 and 12 years tend to be overestimated to be 13 to 16 years old with false positives as high as $0.76$  [^3] NIST FATE Report; higher errors are reported for females than for males, specifically, in 17-year-olds [NIST_FATE_2024]. The performance of facial age estimation algorithms can be affected by several factors like gender, ethnicity, image quality, facial expression, and makeup [anda2020assessing, clapes2018apparent].
@@
Impact of Cosmetic Makeup on Facial Analytics. [ueda2010influence] demonstrated the impact of facial makeup on automatic face recognition. [feng2012quantifying] observed that makeup applied to different facial regions had varying effects on age prediction quantified as Young Index (YI). [eckert2013facial] observed that the Identification Rate (IDR) decreased with increasing amount of makeup.  [chen2013automatic, guo2013face, hu2013makeup] focused on making face recognition more robust to makeup. Authors in [Chen2014ImpactOF] observed that facial cosmetics (both real makeup and synthetically retouched) affect automated age estimation, causing a mean absolute error of up to 5.84 years. They further indicated that lip makeup had the least impact on automated age estimation, followed by eye makeup, followed by full face makeup (eye + lip + foundation). Makeup produces differential effects on apparent (manual) age estimation. 20-year-old women appeared older when wearing full face makeup, while 40- and 50-year-old women appeared younger when wearing makeup [Dayan2015QuantifyingTI, clapes2018apparent, russell2019differential, davis2022your]. Few works discuss the influence of makeup on facial age estimation in minors [anda2020assessing, anda2020deepuage].
@@
Facial Makeup Transfer and Removal. Approaches for makeup transfer relied on GANs for style-transfer networks [li2018beautygan, chen2019beautyglow, jiang2020psgan, nguyen2021lipstick, yang2022elegant, sun2024content]. Later, diffusion models were used [zhang2024stable, sun2024shmt, ruan2025mad]. Makeup removal, on the other hand, poses a more challenging task where the goal is to reveal the face under makeup without adversely affecting the biometric and perceptual features. Anti-makeup [li2018anti] integrated two adversarial networks, both at pixel and feature levels, to generate identity-preserving synthetic non-makeup faces. LADN [ladn_2019] used local disentanglement with adversarial network for makeup transfer and removal.  [zhang2021makeup] replaced convolution layers in GAN with resnet blocks to design makeup removal network. PSGAN++ [PSGAN++_2021] used pose and expression robust spatial-aware GAN for detail-preserving makeup transfer and removal. CLIP2Protect [shamshad2023clip2protect] and DiffAM [sun2024diffam] use adversarial makeup transfer for privacy with capability to generate ‘no makeup’ faces. [ruan2025mad] used domain translation for a wide array of makeup tasks including makeup transfer and removal as a one-for-all diffusion-based framework; but suffers from lack of generalizability, and is limited to only training data and makeup styles. DeBeauty [DeBeauty_2025], is the only reference-free approach to the best of our knowledge that used adversarial de-makeup flow and relocation deformation flow to remove makeup.
III Proposed Method
@@
Our proposed method specifically targets mitigating the impact of makeup on automated facial age estimation and, additionally, facial identity verification. Defending against makeup attacks can prevent minors from bypassing age verification intended for protection against cyber abuse. One strategy may involve collecting data of minors wearing makeup for improving robustness of age estimators; this poses strong privacy concerns. So, we propose a generative model Diffclean, which is designed to mitigate age variations arising due to makeup. There are two challenges in designing this network: (1) we need a robust makeup style generator to synthesize data needed for training the makeup removal model and thus circumvent data collection; (2) we need a robust age estimator to serve as an auxiliary classifier to guide DiffClean to restore age-specific cues. For the makeup style generator, we use EleGANt [yang2022elegant] for locally-editable makeup transfer; see Fig. 2. For the age estimator, we use SSRNet [yang2018ssr] fine-tuned on UTKFace [UTKFace] dataset that has images from different age groups.
@@
Figure 2: Synthetic makeup transfer results produced using EleGANt [yang2022elegant] on example images from UTKFace. Original image (first column), Makeup style reference (second column), and Makeup transferred image (third column).
@@
DiffClean uses the ablated diffusion model (ADM) with guided diffusion from [dhariwal2021diffusion] with improved U-Net architecture containing increased attention heads at multiple resolutions, BigGAN residual stacks for upsampling and downsampling, deeper networks and rescaled residual connections; refer to  [dhariwal2021diffusion] for details. We used text prompts face with makeup and face without makeup for text-guided CLIP directional loss following [sun2024diffam] that detects if makeup is present or not. Makeup removal relies on the combined effect of CLIP-based makeup detection, perceptual utility retention, and restoration of identity and age-specific cues.
@@
Methodology. We describe the terminology and notation here. Original/Clean image (without makeup), $\textbf{I}_{o}$, Makeup image (with real/synthetic makeup), $\textbf{I}_{m}$, and Makeup removed image, $\textbf{I}_{g}$. Consider $\mathcal{M}$ as a real or synthetic makeup application operator, while $\mathcal{M}^{-1}$ represents the makeup removal operator. Thus, $\textbf{I}_{o}\xrightarrow[]{\mathcal{M}}\textbf{I}_{m}$, while $\textbf{I}_{m}\xrightarrow[]{\mathcal{M}^{-1}}\textbf{I}_{g}$. The goal is to model $\mathcal{M}^{-1}$ so that we can recover the clean image, $\textbf{I}_{g}\approx\textbf{I}_{o}$. Our proposed framework DiffClean approximates the role of $\mathcal{M}^{-1}$.
@@
Architecture: The makeup removal framework ($\mathcal{M}^{-1}$) used in this work is a generative model which is a pre-trained ADM [dhariwal2021diffusion] that adds noise to the input image (with makeup), $\textbf{I}_{m}$, to generate the noisy latent image during forward diffusion. The deterministic reverse diffusion through DDIM produces the makeup removed output, $\textbf{I}_{g}$ supervised by CLIP, biometric identity, age and perceptual losses.
Loss functions: We present the loss functions used for supervising the age estimator and the diffusion model.
@@
Age Estimator: SSRNet is fine-tuned using $\mathcal{L}_{WSL}$, a weighted adaptation of self-adjusting smoothed-$\mathcal{L}_{1}$ loss. The adaptive weight ensures stricter penalty for mis-classification in vulnerable age groups [10-29] years as follows:
@@
Refer to [fu2019retinamask] for details about $\mathcal{L}_{SL}$. Note we ablated with different losses such as $\mathcal{L}_{1}$, $\mathcal{L}_{2}$ and Huber loss functions but empirically observed $\mathcal{L}_{WSL}$ loss to perform the best.
DiffClean: During training, we used four losses to fine-tune the diffusion model.
CLIP-based Makeup Loss: The transition from makeup domain to non-makeup domain is guided via textual prompts ‘face without makeup’ and ‘face with makeup’ in CLIP space using ViT-B/16 encoder following [sun2024diffam]. We define the loss as one minus the cosine similarity between the $\mathcal{L}_{2}$-normalized image and text embeddings as follows:
@@
Biometric identity Loss: Our goal involves identity restoration of the makeup removed image with respect to the original image (without makeup). During training with synthetic makeup pairs, we have access to ($\textbf{I}_{o}$, $\textbf{I}_{m}$) to generate $\textbf{I}_{g}$. So, we compute weighted pairwise biometric identity losses as $\mathcal{L}_{og}$ and $\mathcal{L}_{mg}$. This is different than [sun2024diffam] that uses a single identity loss between input and output. The final identity loss is as follows:
@@
Perceptual Loss: To preserve the perceptual features, we use the perceptual loss $\mathcal{L}_{pips}$ with VGG [simonyan2014very] network and $\mathcal{L}_{1}$ loss between generated makeup-removed image $\textbf{I}_{g}$ and input image with makeup $\textbf{I}_{m}$.
Age Loss:
The age loss is computed as the smoothed-$\mathcal{L}_{1}$ loss (introduced in RetinaMask [fu2019retinamask]) between the the ground truth age, $a_{o}$ of the original image, $\textbf{I}_{o}$, and the predicted age, $\hat{a_{g}}$ of the makeup removed image, $\textbf{I}_{g}$. We used the fine-tuned SSRNet model to predict $\hat{a_{g}}$. We used the ground truth age labels as $a_{o}$ provided by [FFHQ_JSON] for FFHQ. We introduce the SSRNet age loss as follows:
$\mathcal{L}_{\text{SSRNet-age}}=\mathcal{L}_{SL}(a_{o},\hat{a}_{g})$ (4)
We further tested with CLIP-based age loss in lieu of SSRNet, where we probed the CLIP model with the generated image, $\textbf{I}_{g}$, using the prompt $\textbf{t}_{age}=\text{{face of }}\{a_{i}\}\text{{-year old}}$, where $a_{i}$ indicates the ground truth age label as follows:
@@
$\begin{split}\mathcal{L}_{total}&=\lambda_{clip}\mathcal{L}_{clip}+\lambda_{id}\mathcal{L}_{id}+\lambda_{lpips}\mathcal{L}_{lpips}+\lambda_{L1}\mathcal{L}_{L1}\\ &+\lambda_{age}\mathcal{L}_{age}\end{split}$ (6)
@@
where loss weights are tuned as $\lambda_{clip}=5.0$, $\lambda_{id}=1.0$, $\lambda_{lpips}=5.0$, $\lambda_{L1}=2.0$, and $\lambda_{age}=0.5$ for SSRNet age loss and $\lambda_{age}=5.0$ for CLIP age loss. Hyperparameters were selected using grid search; see sensitivity analysis in Sec. V-E5. Thus, we design two models: DiffClean with SSRNet age loss ($\mathcal{L}_{age}=\mathcal{L}_{\text{SSRNet-age}}$) and DiffClean with CLIP age loss ($\mathcal{L}_{age}=\mathcal{L}_{\text{CLIP-age}}$).
During testing, our model accepts face images captured under unconstrained conditions, either without makeup or with real or synthetic makeup as input.
Figure 3: Comparison of makeup removal results generated by six baselines and our proposed DiffClean (last two columns) on three examples images from FFHQ dataset. GAN-based baselines (BeautyGAN, LADN, PSGAN++) introduce visual artifacts, while CLIP2Protect alters hair color and style, DiffAM does not effectively remove makeup, and MAD produces distortions on unseen data.
IV Experiments and Analysis
IV-A Datasets